The Linux Distribution HOWTO

by Matt Welsh, mdw@sunsite.unc.edu v2.0, 15 March 1994

This document lists the various software distributions and services available for Linux via mail order, anonymous FTP, and other sources.

Contents

1 Introduction 1 1.1 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 And so on... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Linux Software Distributions 2 2.1 Slackware Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 MCC Interim Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 TAMU Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Linux Support Team Erlangen Distribution (LST) . . . . . . . . . . . . . . . . . . . . . . . . 5 2.5 S.u.S.E. GmbH German Linux CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.6 Debian Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.7 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible . . . . . . . . . . . . . . . . . . 8 2.8 Linux from Nascent CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.9 Trans-Ameritech Linux plus BSD CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.10 The Linux Quarterly CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.11 Unifix 1.02 CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3 Redistributors and Miscellany 14 3.1 Clark Internet Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2 Extent Verlag, LDP Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.3 Fintronic Linux Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4 InfoMagic Linux CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.5 Lasermoon Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.6 Linux Journal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 Linux Systems Labs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.8 Mark Horton Linux Documentation Hardcopy Service . . . . . . . . . . . . . . . . . . . . . . 19 3.9 Sequoia International Motif Development Package . . . . . . . . . . . . . . . . . . . . . . . . 20 3.10 SSC Linux Documentation Project Manuals . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.11 SW Technology Linux Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.12 Takelap Systems Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.13 Unifix Custom Linux CD-ROM Mastering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.14 UPython SLS and Slackware Diskette Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.15 Wizvax Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4 Submissions to this document 24 4.1 New comp.os.linux.announce policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.2 How to submit an entry for this HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4.2.1 Types of submissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.2.2 How to submit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4.3 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

1 Introduction

There is no single distribution of the Linux software. Instead, there are many such distributions, available both via anonymous FTP and mail order on floppy, tape, or CD-ROM. This document is an attempt to present information on many of the available distributions of Linux. This document also contains information on many miscellaneous services and goods available for Linux, ranging from consulting and support to Tshirts and virtual beer.

The purpose of this document is to provide short summaries of the many Linux distributions and mailorder services, and to provide pointers for the reader to find more information. You shouldn’t base your decision of which Linux distribution to use based soley on this document.

This is a PRELIMINARY version of the Distribution HOWTO. It is far from complete; there are many more Linux distributions and services than are listed here. Unfortunately, I have not received submissions from many of the organizations providing these services. Please see Section 4 at the end of this document for information on making a submission. It’s easy and fun and won’t take you five minutes.

1.1 New versions of this document

New versions of this document will be posted periodically (about every month) to the newsgroups comp.os.linux.announce and comp.os.linux.misc. The document is archived on a number of Linux

FTP sites, including sunsite.unc.edu in pub/Linux/docs/HOWTO.

In the future, this document is going to be replaced by a Linux Buyer’s Guide that I am working on. This will contain reviews of each of the major Linux distributions and a list of all distributors. In the meantime, keep those submissions coming.

1.2 And so on...

Please see section “New comp.os.linux.announce Policy” at the end of this document for information on how to make submissions to this document.

Disclaimer: I make absolutely no guarantee as to the correctness of the information, prices, and ordering details given in this document. Several of the entries in this HOWTO are a bit outdated; check the last modified field of each to get an idea. Furthermore, unless otherwise stated the Linux software comes with ABSOLUTELY NO WARRANTY .

I do not endorse or work for any of the distributors listed in this document. I merely include their information here as a service to the Linux community, as a “buyer’s guide”. Inclusion in this document does not mean that I recommend any of the products listed here.

The entries for each distributor are included as they sent them to me. I am not responsible for any spelling errors or mistakes in the content of this document; most of the entries here have only been edited to a limited extent by myself.

Feel free to mail any questions or comments about this HOWTO to Matt Welsh, mdw@sunsite.unc.edu.

And now for the good stuff.

2 Linux Software Distributions

This section describes complete distributions of Linux software, available either via anonymous FTP or from mail order.

In each of the following entries, my own comments are followed by the information from the distributors themselves. You may safely disregard anything I say. :)

An oft-recurring question in the Linux world is, “What distribution should I choose?” This is not an easy question to answer: It depends greatly on your needs and available resources.

The current de facto standard distribution is the Slackware release. Slackware is easy to install and fairly complete. It’s also more or less up-to-date with current versions of Linux software. The Linux Installation HOWTO documents installation of Slackware. I personally recommend Slackware for most new Linux users.

However, if you have a CD-ROM drive, you have more distributions to choose from. I haven’t personally reviewed most of them, and you’re more likely to get help from people on the Net for distributions such as Slackware (which is available freely on the Net, as well as on CD-ROM). On the other hand, some CD-ROM distributions come with documentation and support of their own. All I can say is that a good number of people have had success installing and using the Slackware distribution.

The Debian distribution is currently under development. I have had a chance to play with the beta versions of Debian and it looks very promising. It is being developed and maintained in a semi-closed team of people (instead of by a single individual), so we hope that once Debian matures it will grow into the ”standard” Linux distribution, and the matter will be closed.

The bottom line is that there’s no easy way to pick a distribution out of the air. My own suggestion is to use Slackware. Other people will suggest MCC-Interim, TAMU, and others based on their own experience. By way of disclaimer, I don’t use Slackware myself (in fact, I don’t use any standard Linux distribution), but I have installed it for testing purposes.

2.1 Slackware Linux Distribution

Distributor:

Patrick Volkerding, volkerdi@mhd1.moorhead.msus.edu.

Description:

Slackware Linux is a full featured distribution of the Linux operating system designed for 386/486 computers with a 3.5" floppy. Slackware changes rapidly, but here’s a current (partial) feature list: A series (14 disks):

Base Linux 1.0 kernel OS and utilities. Networking, UUCP, gcc/g++ 2.5.8, libc 4.5.24, and more. E series (5 disks):

Emacs 19.22. F series (1 disk):

FAQs, HOWTOs, and other documentation. OI series (3 disks):

ObjectBuilder 2.0 for X. X series (5 disks):

Base XFree86 2.1 X windows system with fvwm. XAP series (2 disks):

X windows applications such as seyon and ghostview. XD series (3 disks):

X windows program/server development. XV series (2 disks):

XView 3.2 release 5, Open Look Window Manager. Y series (1 disk):

Games from BSD, such as ”hunt”.

Availability:

The home site is ftp.cdrom.com, where the latest distribution can be found in /pub/linux/slackware. To make it easy to download, the disks can also be found pre-zooed in /pub/linux/zooed slackware. (Zoo is just a compression utility for MS-DOS).

Ordering:

FTP only, although various independent distributors provide it on disk, floppy, and CD.

Entry last modified:

5 Nov 93.

2.2 MCC Interim Linux

The current MCC-Interim release is based on 0.99.pl10, which is quite robust. If you are one of those people who demands to be on top of the current release, you’ll need to upgrade your kernel (easy to do) and possibly your version of GCC and libraries. However, I see no reason why 0.99.pl10 shouldn’t be good enough, at least to get you started.

Distributor:

Dr. A. V. Le Blanc, LeBlanc@mcc.ac.uk.

Description:

Base Linux installation. Complete sources and patches for all included software are available.

Full details vary from release to release; see /pub/linux/mcc-interim/*/Acknowledgements at

ftp.mcc.ac.uk. Roughly: this includes the kernel (with source), C, C++, groff, man pages, basic utilities, networking. The binaries fit on 6 or 7 floppies. It is also possible to have them on a DOS or Linux partition on a hard drive (plus one floppy) or on an NFS-mountable partition (plus two floppies). New versions appear at 2-3 month intervals, depending on various factors.

Availability:

By anonymous ftp from ftp.mcc.ac.uk; mirrored at tsx-11.mit.edu, nic.funet.fi, and elsewhere.

At ftp.mcc.ac.uk, in /pub/linux/mcc-interim.

Miscellaneous:

Suggestions and contributions are welcome.

Entry last modified:

31 Oct 93.

2.3 TAMU Linux Distribution

Distributor:

Dave Safford, Texas A&M University, dave.safford@net.tamu.edu.

Description:

TAMU.99p12+ is the latest release in the TAMU linux series. Unlike previous releases, this one includes both integrated source and binary sets, with the entire binary set created from a single top level source make. This ensures that all programs are compiled and linked with the same current tools and libraries, and guarantees availability of working source for every program in the binary set. In addition, the new boot diskette fully automates the installation process, including partitioning, lilo bootstrapping, and network configuration. Installation requires no rebooting, and requires the user to know only the host’s name and IP address. At every step of installation, the program provides intelligent defaults, making it a snap for novices, while allowing experts full flexibility in setting installation parameters. Reliability has been improved over past TAMU installation by the use of labels on all disk images, so that the program can detect and recover from bad or misordered diskettes. This release is a full featured package, including XFree86-1.3, emacs-19.18, net-2, bootutils, and sources for all installation programs (without any use restrictions :-).

Availability:

The latest TAMU release is available by anonymous ftp from net.tamu.edu:pub/linux.

Entry last modified:

31 Oct 93.

2.4 Linux Support Team Erlangen Distribution (LST)

Distributor:

The Linux Support Team Erlangen, a small group of students at the University of Erlangen-Nuernberg. Contact Stefan Probst (snprobst@cip.informatik.uni-erlangen.de) or Ralf Flaxa (rfflaxa@informatik.uni-erlangen.de). Probably, there will be a collective address soon.

Description:

The LST distribution’s goal is to provide a solid, reliable, easy to install (even for beginners) and well-documented system. We are not hunting for the newest kernel or gcc versions. We do updates when they are necessary or provide really new functionality, are well tested, integrated in the system and working smoothly with the rest of the system. The distribution consists of a base system and additional packages. Currently the following packages are available: doku (doc), text, tex, develop, xdevelop, xbasis, xappl, xemacs, tinyx, network, grafik, src, misc. The complete system is 50 HD disks and 1500 pages printed documentation including the LDP guides (IGS,KHG,NAG), HOWTOs, FAQs, the german Linuxhandbuch, install-guide and many other useful documents we collected over the time. Our distribution is preconfigured for german users and comes with a 50 page step by step installation guide that leads you through the menu-driven installation scripts (both in german). We started this distribution to help newcomers with their first steps into Linux. Therefore our scripts are smart enough to handle all of the “dirty work” like setting up system configuraton (including LILO, modem, mouse, mounts, mtools, access to DOS, users, X11) and network configuration (TCP/IP, Routing, Mail, News, UUCP, SLIP). Historically (and of course easier for us) most of the documentation for this distribution is written in german, sorry, but we are planning to translate those documents in english, but at the moment we don’t have time for this. Any volunteers are welcome!

Our scripts also accept to install SLS and Slackware packages, but with no warranty if they work well together with the rest of the system.

Availability:

Via anonymous ftp: ftp.uni-erlangen.de under pub/Linux/LST.Distribution, or on 3.5" disks

(also on one QIC-80 tape).

Ordering:

order at following adress: Stefan Probst, In der Reuth 200, 91056 Erlangen, Germany.

• complete documentation (IGS, LHB, KHG, GDB, NAG, HOWTOs, FAQs, Installguide) together over 1500 pages for DM 139,– (plus shipping).

• complete documentation plus complete distribution on QIC-80 tape (about 65 MB) for DM 199,– (including shipping in Germany).

• complete documentation plus complete distribution on about 50 3.5" disks for DM 269,– (including shipping in Germany).

Entry last modified:

21 Dec 93.

2.5 S.u.S.E. GmbH German Linux CD-ROM

Distributor:

     S.u.S.E. GmbH
     Gebhardtstr.  2
     90762 Fuerth
     Germany

Description:

This is a CD especially for German users. It contains a german version of the Slackware Distribution and two additional german distributions based on Slackware and SLS. It contains additional software, such as the Postgres system, the pbm-Tools and lots of other tools and utilities (ctwm window manager, GREAT environment, andrew toolkit, the POV-ray raytracer and so on). The source code of the complete system is included, too, as is a live system to enable using software without the need to install everything to the harddrive. The current version of the CD contains version 1.2.0 of the Slackware distribution containing version 1.0 of the linux kernel, version 2.5.8 of the GNU-Compiler and XFree86[tm] 2.1. A german installation guide will help new users to successfully install the system. The CD ist updated every three months, so the next one will be released in July 1994.

Ordering:

You can order by mail, by phone (49-911-74053-31), by Fax (49-911-7417755) or by sending email to

bestellung@suse.de.

Price: The price for a single CD is 89,– DM for new customers. If you are already a S.u.S.E.-customer, you will get the CD for an update-price of 78,– DM. There is also the posibility to subscribe for one year; then you will receive 4 CDs for the price of 230,– DM. Disks: Of course, you can get the german version of the Slackware Distribution on 3,5" disks. You can get either a basis version of the distribution (about 28 disks) for 89,– DM, the ”standard” version (about 38 disks) for 119,– DM or the ”full” version (about 59 disks) for 159,– DM. The update prices are 69,– DM, 94,– DM and 139,– DM respectively. Documentation: S.u.S.E. is selling high-quality printed issues of the books from the Linux Documentation project (LDP) for people that would like to have a real book and not only printed loose papers. The prices are 24,– DM for Matt Welsh’s Linux Installation and Getting Started, 39,– DM for Olaf Kirch’s Network Administrators’ Guide and 29,– DM for Michael K. Johnson’s Kernel Hackers’ guide. Additionally, S.u.S.E. offers the complete O’Reilly series which is the best documentation available for Unix in general. Of course you can order the well known german Linux books: Deutsches Anwenderhandbuch for 49,– DM and Linux - vom PC zur Workstation for 38,– DM. Motif: Metrolink Motif for Linux is available for the price of 169,– DM for the runtime system only and for the price of 288,– DM you will get the runtime and development system.

Miscellaneous:

Service and support for our customers by Email, Fax, snail mail and phone (hotline Monday and Thursday from 13:00h-17:00h.) Free info material and catalogues of CDROMs and books. (All Infomagic Unix CDs available). Selling commercial software for Linux. Development of individual software for Linux and other Unix systems. Almost anything concerning Linux - just ask!

Entry last modified:

23 May 1994

2.6 Debian Linux Distribution

Distributor:

Ian A. Murdock, imurdock@gnu.ai.mit.edu.

Description:

For those of you who are not familiar with Debian, it is an effort to create a well thought-out, powerful, flexible yet complete Linux distribution. The motivations behind Debian are detailed in the Debian Manifesto, a copy of which can be acquired as described below. For up-to-date information, please look at the files in the directory

/pub/Linux/distributions/debian/info at sunsite.unc.edu. If you do not have access to the

Internet or FTP, you may obtain printed copies by sending a self-addressed, stamped envelope to:

The Debian Linux Association
Station 11
P.O. Box 3121
West Lafayette, IN  47906
USA

This information includes how Debian can be obtained (via FTP or mail-order), why Debian is being constructed (the Manifesto) and other general information (including how to join the Debian mailing lists).

Availability:

Debian has not yet been ”officially” released, but BETA releases are available to the general public at

sunsite.unc.edu in the directory /pub/Linux/distributions/debian.

Entry last modified:

21 Feb 94.

2.7 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible

Distributor:

     Yggdrasil Computing, Incorporated
     4880 Stevens Creek Blvd., Suite 205
     San Jose, CA 95129-1034
     toll free (800) 261-6630, (408) 261-6630, fax (408) 261-6631
     info@yggdrasil.com

Description:

Yggdrasil Plug-and-Play Linux is a complete CD-ROM distribution of the Linux operating system. It includes a great deal of software—nearly every package that you would expect to find on a complete UNIX system is available. A complete file list is available via FTP from yggdrasil.com. The Linux Bible is a compendium of Linux documentation, including 3 books from the Linux Documentation Project, The Yggdrasil installation manual, and the complete set of Linux HOWTO guides. At the top of the sophistication hierarchy, programmers who want to explore or add the occasional feature know that Yggdrasil is the Linux distribution with a fully buildable source tree and with the ability to automatically trace installed files back to their sources. Users who want maximum performance will appreciate that the major system components have been recompiled with -O6 optimization, and the SCSI clustering which reduces the build time on the source tree from 28 to 22 hours (over 20%) on 486DX2-66. Using IDE? Activate the multisector IDE code! Everybody, especially new users, will appreciate the Plug-and-Play operation for which Yggdrasil’s product is named. Put the media in a computer with supported hardware, turn the computer on, and it’s running everything, straight from the CDROM. The login screen lists a number of preconfigured user names, including ”install”, which installs the system, giving paragraphs of explanation about every question that it asks the user. The install script even searches for a modem, and, upon finding it, configures mail and UUCP so that mail sent to an internet address is transparently delivered through a bulletin board system at Yggdrasil. X windows configuration is automated too, prompting the user for configuration information the first time xinit is run. From X windows, a graphical control panel allows simple “fill in the blanks” configuration of networking, SLIP, outgoing UUCP, the printer, NNTP, and many other features that previously required the knowledge of a system administrator to configure. For more information, send mail to info@yggdrasil.com, FTP to yggdrasil.com, or contact us by any convenient method.

Ordering:

Plug-and-Play Linux costs $39.95 is and available directly from Yggdrasil or from your local computer, software or technical book store. If Plug-and-Play Linux is not available from your favorite reseller, help promote Linux by making it your mission to change that. Give your reseller our phone number and demand that they carry Plug-and-Play Linux. Yggdrasil offers a $10 discount for upgrades or crossgrades. Send us your old Yggdrasil release and a check for $29.95 + $5 shipping and handling to upgrade. Or do the same, but send us a competing distribution such any version of SCO, Esix, minix, or one of the CD’s with the slackware floppy images, and tell us where you got it, so that we can make sure that your favorite reseller carries our products too.

Miscellaneous:

Yggdrasil also sells OSF/Motif and The Linux Bible, a compilation of works from the Linux Documentation project. In addition to a copy of the Plug-and-Play Linux manual, The Linux Bible includes Linux Installation and Getting Started, Network Administrator’s Guide, Kernel Hacker’s Guide, and HOWTO guides on hardware compatibility, distributions, dos emulation, ethernet, floppy tape, installation, mail, networking version 2, MGR (an alternative to X windows), X windows, usenet news, printing, SCSI, serial communications, sound, and UUCP. The Linux Bible costs $39.95 and is printed on recycled paper. $1 from every copy sold is donated to the Linux Documentation Project. OSF/Motif costs $149.95, $5 of which is donated to the development of a free Motif clone.

Entry last modified:

16 May 1994

2.8 Linux from Nascent CD-ROM

Distributor:

Nascent Technology

Description:

The Linux from Nascent CDROM is a new distribution of the Linux operating system which includes over 400 mbytes of source code, binaries, and documentation for Linux and applications. It features automated root, swap, package, network, and user account installation from CDROM. Linux can be can be run directly from the CDROM and floppy. The Nascent CDROM features Xwindows, Openlook, TeX, GNU compiler and utilities, Magic and Spice electronic design tools, and over 100 high resolution images translated from Kodak PhotoCD(TM). Each source archive is distributed with an associated notes file to allow you to browse and install applications using a consistent interface. A listing of the contents of the Nascent CDROM as well as a current copy of the CDROM announcement and order form may be obtained via anonymous ftp at netcom.com:/pub/nascent.

Ordering:

The Linux from Nascent CDROM, Version 1.0, is only $39.95 plus shipping and handling. Nascent also offers the Linux from Nascent Plus package for only $89.95, which includes six months of email support and a 30 the CDROM with your CDROM purchase. Nascent accepts Mastercard, VISA, checks, and money orders as payment. To order your Linux from Nascent CDROM, mail, email or fax a completed order form to:

Nascent Technology
Linux from Nascent CDROM
P.O. Box 60669
Sunnyvale CA 94088-0669 USA
Tel: (408) 737-9500
Fax: (408) 241-9390
Email: nascent@netcom.com

Entry last modified:

28 Nov 93.

2.9 Trans-Ameritech Linux plus BSD CD-ROM

Distributor:

Trans-Ameritech corporation.

Description:

CD-ROM based on the Slackware distribution of Linux with all the sources plus full uncompressed filesystem plus NetBSD source and binary distribution. The current release is “Spring 94”. The two previous releases cover Linux from 0.99 p.9 (on the SLS release) to 0.99 p.14 and have FreeBSD and the original 386BSD.

• The installation is highly automated by the Slackware scripts fine tuned to this CDROM but leaves a lot of room for customization.

• In addition to standard device drivers in Linux kernel, BIOS-less aha1522 SCSI is supported (it covers Sound Blaster 16 SCSI too).

• To minimize the possibility of hardware conflicts many extra kernels are provided for different configurations. They are usable for installation and normal use.

• Many on-line documents are provided for quick reference, including the Linux Documentation Project files in source, dvi and ps formats.

• A lot of applications are included:

– The C/C++ compiler GNU GCC 2.5.8,

– GNU and international versions of the ispell spell-checker.

– The communications apps: term 1.1.4, minicom, Seyon (X-Windows based).

– Editors: elvis (vi clone), joe, jove

– PostScript clone ghostscript 2.6.1,

– Network package with news and email:

– TCP/IP (net2debugged), UUCP, SLIP, CSLIP

– Object oriented GNU Smalltalk 1.1.1, and the Smalltalk Interface to X. (STIX)

– TCL/Tk (Powerful scripting language with Motif-like X interface)

– SPICE for electric engineering

– Several window managers - openwin, twm, gwm, fvwm.

– Easy X-Windows configuration with many real-life examples.

– Many X applications

– Interviews libraries, include files, and the doc word processor and idraw drawing program.

– Typesetting: TeX, LaTeX, xdvi, dvips, Metafont, groff

– Ingress and Postgress databases.

– Multimedia: mpeg video and sound applications. Mosaic 2.2. xgopher.1.3.2, tracker,

  adagio04, speak-1.0.

– Gif pictures and sounds in various formats.

– Andrew multimedia word processor with hyperlinks.

– FlexFax send and receive fax on either class 1 or class 2 fax-modems.

– A lot more.

• To help the first time Linux users many documentation files are provided that are readable from DOS even before installing Linux.

• All the sources are available on the CDROM. The most often needed sources are uncompressed and can be used directly from CDROM.

• An uncompressed Linux filesystem is available for reference and disk space conservation. You can run programs directly from CDROM! There is a large info directory for on-line reference and many manpages.

• For hacker’s reference an uncompressed FreeBSD source tree is provided.

Availability and Ordering:

You can order by e-mail, by phone (408)727-3883 or fax (408)727-3882. (Email:

Roman@Trans-Ameritech.com)

If you prefer to send a cheque, our address is:

     Trans-Ameritech Enterprises, Inc.
     2342A Walsh Ave
     Santa Clara, CA 95051
     USA

If you order with a credit card (VISA, MC, American Express) please indicate the card number, expiration date and your mailing address. The order will be processed and the CD shipped the same day. The price for a single CD-ROM is $30 if you are an internet user and found the info on the net (you must have an e-mail address). Shipping and handling in US is $5. Overseas is $8. COD is available in the US only for $4.50. California residents please add sales tax. Annual subscriptions (4 releases) are available for $80 plus S&H. (note: there are 4 shipments in a subscription) Example subscription in US is: $80 + $5 x 4 = $100 Subscription in Europe/Japan etc. is: $80 + $8 x 4 = $112

Miscellaneous:

We answer technical questions about our distribution by e-mail sent to: roman@trans-ameritech.com within 24 hours! Trans-Ameritech is a well established company with efficient organization to guarantee on-time printing and shipping of this product. If you have a 3.5 inch boot floppy and have ftp access,

please look on sunsite.unc.edu in /pub/Linux/distributions/TransAmer for a new boot floppy

image with support for many CDROM drives.

Entry last modified:

19 April 1994.

2.10 The Linux Quarterly CD-ROM

Distributor:

Morse Telecommunication, Inc.

Description:

The Linux Quarterly CDROM contains the complete contents of tsx-11.mit.edu, the one of the most popular Internet Linux sites which provides both source and binaries of major Linux distributions, utilities, source code and documentation. This includes Slackware, SLS, MCC and Debian releases of Linux. Additionally, beginning with the Spring 1994 edition, the complete contents of prep.ai.mit.edu, the repository of the FSF’s GNU source archives, is included. For first-time users, this CDROM contains a Microsoft Windows(tm) front-end to assist the user in creating boot disks that can be used to install Linux directly from The Linux Quarterly CDROM. Support for the UMSDOS filesystem is now also This gives the user the ability to install Linux directly to an MSDOS system without having to repartion their hard drive. Ninety days of technical support is provided with each disc at no additional charge.

Availability:

     Morse Telecommunication, Inc.
     26 East Park Avenue, Suite 240
     Long Beach, NY 11561
     Orders:  (800) 60-MORSE
     Tech Support:  (516) 889-8610
     Fax:  (516) 889-8665
     Email Orders:  Order@morse.net
     Email info:  Linux@morse.net
     Hours:  9AM - 5PM EST, Monday through Friday.  Fax available 24 hours.

Ordering:

Orders may be placed via phone, fax or email as listed above. The Linux Quarterly CDROM is priced at $29.95 plus $5 shipping and handling per order. Existing customers may upgrade for $22.95 plush S/H. All orders generally ship the same day if received by 4PM EST.

Miscellaneous:

The contents of the CDROM, the cover artwork in JPEG format, and the text of the USENET announcement can be found on the following FTP sites:

     tsx-11.mit.edu:/pub/linux/advertisements/TLQ-Spring94.tar.z
     sunsite.unc.edu:/pub/Linux/distributions/cdrom/TLQ-Spring94.tar.z

Entry last modified:

23 May 1994

2.11 Unifix 1.02 CD-ROM

Distributor:

Unifix Software GmbH, Braunschweig, Germany

Description:

This is a Linux CD-ROM distribution with emphasis on easy and fast installation. Though it is possible to install everything on the harddisk, we do not recommend so, because it runs fast enough directly from CD. For example starting emacs the first time (in text mode) takes:

  From single speed Mitsumi: 24 s
  From double speed Mitsumi: 11 s
  From Toshiba 3401:          7 s

The dynamic buffer cache of Linux takes care of the second and subsequent times a program is started, which needs less than one second. Our system requires about 5 MB on the hard disk for configuration files and system management tools. Additionally we recommend a 16 MB swap partition. The distribution contains most of the standard programs in current versions, e.g. Linux 0.99.15e, XFree 2.0, emacs 19.22 and gcc 2.5.7. Full preconfigured source for everything is included.

Because Unifix is a european distribution it provides full support for iso8859-1 character sets. From the shells and emacs through ls, TeX and the print system everything supports 8-bit characters. Printing is supported through System V compatible printer drivers which can be controlled through lp’s -o options. Printer drivers are included for dumb text mode printers, for postscript- or ghostscriptcompatible printers and for networked printing. These drivers know how to guess the type of the files so e.g. compressed man pages or DVI files can be printed directly.

Ordering:

Unifix is available only on CD. It comes in a Unifix/Linux binder with two boot floppies and about 70 pages of installation instructions in german. An english version will be available Apr 94. The price is 159 DM (about $100) and includes taxes and shipping; Eurocard/Mastercard/Visa are accepted.

Telephone       +49 (0)531 515161
Fax             +49 (0)531 515162
Mail            Unifix Software GmbH
                Postfach 4918
                D-38039 Braunschweig
                Germany

Entry last modified:

18 Feb 94.

3 Redistributors and Miscellany

This section lists re-sellers or re-distributors of the Linux distributions listed above. In other words, the people selling the below software more than likely do NOT maintain or support the software itself.

This section also lists distributors selling commercial software for Linux (such as Motif) and miscellany, such as documentation.

3.1 Clark Internet Services

Distributor:

Clark Internet Services, C/O Stephen Balbach

Description:

The latest version of Linux on disk and tape direct from the net at affordable prices. 120MB installed on 3.5" disks. 90MB installed on 5.25" disks.

Ordering:

Slackware on 32 3.5" disks: $50 (USPS 2nd day). SLS on 32 5.25" disks: $40 (USPS 2nd day). Installation and Getting Started (c) Copyright Matt Welsh. 150+page laser printed, professional bound manual. Everything to get installed and running in one easy book. $15 if ordering set (at cost price!). $20 seperate.

Availability:

For more complete information, including a multipage description of Linux, a hardware compatibility sheet, and more detailed package descriptions send mail to linux-all@clark.net for auto-reply info (30k of text).

Ordering:

Check, money-order

Stephen Balbach
5437 Enberend Terrace
Columbia, MD 21045

Credit Card orders call 410-740-1157 (Visa, MasterCard, AmXpres)

Entry last modified:

31 Oct 93.

3.2 Extent Verlag, LDP Distribution

Distributor:

Extent Verlag Berlin, Germany

Description:

Extent has published the Linux Installation and Getting Started manual, Version 2.0 by Matt Welsh in order to make it available to Linux users within Germany and Europe. Inspite of being non-commercial, the manual was printed with 2540 dpi and bound in soft cover. It’s handy paperback format is 148 x 210 mm. Other LDP manuals are being prepared as soon as possible. Extent also tries to make SLACKWARE available to everybody within Germany and Europe just for the ordinary diskette price. This means, you pay only the diskettes and get the newest SLACKWARE distribution for free, just in the sense of Free Software. The 3.5" diskettes are fully error-checked, of course.

Availability:

Extent Verlag Berlin, Postfach 12 66 48, D-10594 Berlin, Germany. phone +49 30 3244021, fax +49 30 3249685.

Ordering:

Linux Installation and Getting Started, 192 pages. ISBN 3-926671-12-2. Costs within Germany DM 15.80 (tax and shipping included), within Europe DM 16.50 (shipping included). Full Slackware (latest version) on 50 diskettes. Within Germany, DM 100.00 plus DM 10.00 for shipping. Within Europe, DM 100.00 plus DM 20.00 for shipping. Ordering by sending cheque to address above or money order (don’t forget your address!) to Extent Verlag Berlin, Germany Postbank Berlin, BLZ 10010010, bank account 1769-104 or every book store.

Entry last modified:

25 March 1994.

3.3 Fintronic Linux Systems

Distributor:

Fintronic Linux Systems

Description:

We sell fully installed, custom configured Linux systems for about the price which you’d pay for the hardware alone. We offer desktop and notebook machines. We ship worldwide and accept payment by check and credit cards. If you have any questions or would like to be added to our mailing list, send

mail to linux@fintronic.com.

Availability:

For  our  latest  pricelist,

finger linux@fintronic.com or Web to http://www.fintronic.com/linux/catalog.html. Prices

change frequently as we are constantly looking for the best deals for our customers.

Ordering:

email:  linux@fintronic.com
  fax:   +1.415.325-4908
voice:   +1.415.325-4474
 Mail:   Fintronic USA, Inc.
         1360 Willow Rd., Suite 205
         Menlo Park, CA 94025
         USA

Entry last modified:

24 Mar 94.

3.4 InfoMagic Linux CD-ROM

Distributor:

InfoMagic, Inc.

Description:

Complete snapshot of tsx-11 and sunsite.unc.edu linux archives taken at 2 month intervals (Dec, Feb, Apr, ...). The disc is in ISO-9660 format with Rock Ridge Extensions. A number of “distributions” are included: SLS, Slackware, MCC, and TAMU. Versions are as available at the date of capture.

Availability:

Readme and index files are available via anonymous ftp from ftp.uu.net:/vendor/InfoMagic/cds. Other info is available via email (info@InfoMagic.com), FAX (609-683-5502) or Tel (800-800-6613 or 609-683-5501).

Ordering:

The disc is priced at $15/copy. Shipping within the US is $5, outside $10. Subscriptions are available for $75/year with shipping of $25 in the US and $35 outside. Orders may be placed by email, FAX or phone. We accept Mastercard and Visa, and will begin accepting American Express after 1/1/94.

Entry ast modifiied:

22 Dec 93.

3.5 Lasermoon Ltd.

Distributor:

Lasermoon Ltd., info@lasermoon.co.uk, support@lasermoon.co.uk.

Description:

The following Linux products are distributed and supported, (although there are many more of interest):

• The Yggdrasil LGX CD-ROM

• The Infomagic Linux Developer Resource CD-ROM

• The Linux Journal (Monthly Mag), European Distributors Availability:

Lasermoon Ltd, 2a Beaconsfield Road, Fareham, Hants, England. PO16 0QB. Voice +44 (0) 329 826444. Fax +44 (0) 329 825936. Email: info@lasermoon.co.uk (General enquiries),

support@lasermoon.co.uk (Support desk), lj@lasermoon.co.uk (Linux Journal desk).

Ordering:

LGX currently 44.95ukp, LDR currently 12.95ukp All prices exclude postage and VAT at the UK rate (17.5 Discounts available - please call. Visa, Mastercard, Access, EuroCard accepted.

Miscellaneous:

We provide snapshots of any part of any product on (almost) any media. Call us for details. Free catalogue of many other CD-ROM Freeware products for UNIX, DOS and Novell available on request (either by email or post). Books by SSC and O’Reilly Associates (10

Entry last modified:

16 Feb 94.

3.6 Linux Journal

Publisher:

Linux Journal, P.O. Box 85867, Seattle, WA 98145-1867. Phone (206) 527-3385 (subscriptions) or (206) 524-8338 (advertising). FAX (206) 527-2806 (subscriptons) or (206) 526-0803 (advertising).

Description:

Linux Journal is a monthly publication covering the Linux Community. Most material in LJ is new (not reprinted from Usenet). Each issue includes columns and articles on Linux programming, Free Software Foundation issues, systems administration, Questions and Answers, interviews and more. LJ is a professional-quality magazine for the Linux community.

Availability:

Subscriptions are $19/year (US), 24/year(Canada/Mexico),29/year (elsewhere).

Ordering:

Payment can be made using VISA, MasterCard or American Express (make sure you include the credit card number, expiration date and signature). We can also accept checks in native currency. If you have a question concerning appropriate method of payment, phone or FAX our subscription numbers or send e-mail to subs@ssc.com For security reasons we discourage sending credit card numbers via e-mail.

Miscellaneous:

If you are interested in advertising in LJ, contact Joanne Wagner by phone or e-mail at joanne@fylz.com. Article queries, new product announcements and other editorial material should be sent to our address above or e-mailed to ljeditor@sunsite.unc.edu. Generic questions can be

sent to linux@fylz.com.

Entry last modified:

26 May 94.

3.7 Linux Systems Labs

Distributor:

Linux Systems Labs, dirvin@vela.acs.oakland.edu.

Description:

Slackware (all 50 disks and I&GS) for $69.95(updated weekly), resell the Yggdrasil CDROM, Motif by Metrolink for $175.00 and bundle any version of Linux with Motif (version 1.2.4) for $215.00. We currently print LDP documents on a 600 DPI Duplexed on a laser printer. We have begun publishing The Linux Bible: The GNU Testament, Spring Quarterly Edition, which includes Linux Installation and Getting Started, the Yggdrasil Plug-and-Play Linux Manual: Summer 1994, The Linux Network Administrators’ Guide by Olaf Kirch, The Kernel Hackers’ Guide by Michael K. Johnson, and the following HOWTOS: distribution, dosemu, ethernet, ftape, installation, mail, mgr, net-2, news, printing, scsi, serial, sound, uucp, xfree86. Also we just started selling a commercial database called /rdb for Linux by Revolutionary Software for $149.00. (It’s quite impressive.) Contact us for more information.

Availability:

Linux Systems Labs, 18300 Tara Drive, Clinton Twp MI 48036. Phone (313)954-2829, (800)432-0556, fax (313)954-2806.

Ordering:

See above for prices. We take MC and Visa and will discount all products (except /rdb) 20% for internet customers.

Entry last modified:

11 May 1994.

3.8 Mark Horton Linux Documentation Hardcopy Service

Distributor:

Mark Horton Associates, mah@ka4ybr.atl.ga.us. Description:

The Linux Hardcopy Service provides printed, punched, and bound copies of the various Linux LDP manuals, FAQs, HOWTOs and other related publications. The goal is to provide both new and experienced Linux users with laser printed, high quality documentation. This is particularly useful to new users without access to ftp or the facilities required to print .dvi format, TeX, or PostScript files.

Availability:

Mark Horton Associates; P.O. Box 747; Decatur, GA US; 30031. Phone: 1.404.371.0291, e-mail:

mah@ka4ybr.atl.ga.us.

Ordering:

(Please e-mail mah@ka4ybr.atl.ga.us for complete info, only prices are given here.)

1. Linux Installation and Getting Started, 150 pages, $20.00

2. Linux Network Administrators’ Guide, 250 pages, $30.00

3. Linux Kernel Hackers’ Guide, 120 pages, $20.00

4. Das LinuXHandBuch, 250 pages, $30.00

5. LILO Technical Overview and User’s Guide, 35 pages, $10.00

6. FAQs and HOWTOS, 250 pages, $30.00

7. Shells, shells, shells (man pages), 100 pages, $15.00 8. Any other documentation you wish printed... make a request, I’ll print a copy, look it over, and

provide a price... if it’s of general interest I’ll offer additional copies also. (Somehow I feel like I just committed to ending up with a LOT of printed documentation on my shelves!)

I can’t take credit cards (you wouldn’t believe what the bank wants to process those things!) Personal checks, money orders, cash are all ok. For very large orders from businesses and educational sites I guess I could take purchase orders. Please add $2.00 per manual for shipping (ups ground - add an additional $10.00 for ups blue label). Special overnight shipping can be arranged upon request. Call for quantity discount pricing. $1.00 per manual sale goes to the Free Software Foundation. $1.00 per manual sale goes to Linus and the LDP’s virtual beer fund. Please make checks payable to Mark Horton Associates, address given above.

Miscellaneous:

Custom t-shirts available on request—send e-mail.

Entry last modified:

15 January 1994

3.9 Sequoia International Motif Development Package

Distributor:

Sequoia International, Inc., 600 West Hillsboro Blvd, Suite 300, Deerfield Beach, FL 33441, Tel: (305) 480-6118, FAX: (305) 480-6198, info@seq.com

Description:

Sequoia International, Inc. has a complete Motif 1.2.3 Runtime and Development package called SWiM 1.2.3 available for $149.95. In addition to providing shared library versions of libXm and libMrm, the following is included in each package: The Window Manager (mwm), Shared Libraries (libXm, libMrm), Static Libraries (libXm, libMrm, libUil), UIL Compiler, Header and Include Files, Complete On-Line Manual Pages, Source Code to OSF/Motif Demo Programs, Complete OSF/Motif Users Guide.

Requirements:

Linux 0.99pl13 or higher, libc 4.4.4, XFree 2.0, 12M Free Disk, 8-12M RAM Suggested.

Ordering:

USA: Sequoia International, Inc. (305-480-6118), info@seq.com. Japan: Fortune Co., Ltd (03-5481-8974). England: Lasermoon Ltd. (+44-0-329-826444). Australia: Space Age Import-Export Proprietary (61-7-266-3418).

Entry last modified:

24 March 1994.

3.10 SSC Linux Documentation Project Manuals

Distributor:

SSC, Inc., sales@ssc.com.

Availability:

SSC has printed and comb bound the Linux Installation and Getting Started manual, Version 2 to make it available for those who do not have the capability to print it themself and to support our customers who buy Linux (Yggdrasil) from us. It is printed double-sided with a cover.

Ordering:

The manual is $15 plus shipping ($3 in the U.S.). We can accept credit card orders (Visa, MasterCard or AmEx). Orders can be phoned in (206-FOR-UNIX/206-527-3385), FAXed (206-527-2806) or mailed to SSC, P.O. Box 55549, Seattle, WA 98155.

Miscellaneous:

SSC also publishes a series of reference cards on Unix and Unix-related programs such as Emacs, VI, Korn Shell, C language, etc. We also sell the Yggdrasil Linux distribution and Free Software Foundation books. Call or e-mail for a free catalog.

Entry last modified:

8 Feb 94.

3.11 SW Technology Linux Systems

Distributor:

SW Technology

Description:

SW Technology has been selling Linux workstations locally with high customer satisfaction. Our installation of Linux follows the style of the Slackware distribution, with system specific kernel image and configuration files, to ensure optimal performance. The installation is a fairly complete distribution of available Linux utilities, including GCC, X, Interviews, Xview, Tcl/Tk, Emacs, TeX/LaTex, Groff, etc. We custom configure systems: our clients will get the hardware components and software configuration per their specification.

Availability:

For upto date prices/info, please

email/finger      swt@netcom.com
anonymous ftp     netcom8.netcom.com:/pub/swt/info
phone             (214) 907-0871
mail              SW Technology
                  251 West Renner Suite 229
                  Richardson, TX 75080

Entry last modified:

15 Jan 94.

3.12 Takelap Systems Ltd.

Distributor:

Takelap Systems Ltd., info@ddrive.demon.co.uk.

Description:

The SLS 1.04 distribution on diskettes. 25 3.5" or 30 5.25" diskettes including 0.99.13 kernel, X11R5, TeX, doc (WYSIWYG), GNU s/w development tools and more. The SLS 1.04 Linux Distribution on CDROM: 0.99.13 kernel, X11R5, TeX, Andrew, GNU s/w development tools. May be installed and optionally run from CDROM mounted remotely. Many installation and operational options. The Yggdrasil LGX Fall 93 Linux distribution on CDROM: 0.99.13 kernel, X11R5, TeX, Andrew 5.1, Postgres 4.1, GNU s/w development tools.

Availability:

Takelap Systems Ltd., The Reddings, Court Robin Lane, Llangwm, Usk, Gwent, United Kingdom NP5 1ET. tel: +44 (0)291 650357, fax: +44 (0)291 650500, email: info@ddrive.demon.co.uk.

Ordering:

SLS on Diskettes, price ukp42.00 +ukp1.50 p&p + VAT. SLS CD-ROM, ukp66.00 +ukp1.50 p&p + VAT. Yggdrasil LGX CD-ROM, ukp40.00 +ukp1.50 p&p + VAT. Visa and Mastercard accepted.

Miscellaneous:

Free catalogue of CDROMs and books (mostly UNIX related) on request. Entry last modified:

31 Oct 93.

3.13 Unifix Custom Linux CD-ROM Mastering

Distributor:

Unifix Software GmbH, Braunschweig, Germany Description:

We use our own (Linux based) premastering system to make custom CDs which include the customer’s programs in addition to our Linux distribution. This enables our customers to sell their Unix based applications together with operating system tested and ready to run. Our CD recorder makes it possible to create prototypes and very small series at interesting prices.

Availablility:

Ask us for more information or prices:

Telephone       +49 (0)531 515161
Fax             +49 (0)531 515162
Mail            Unifix Software GmbH
                Postfach 4918
                D-38039 Braunschweig
                Germany

3.14 UPython SLS and Slackware Diskette Labels

Distributor:

UPython Computer and Network Services

Description:

Labels for Softlanding and slackware disks

Availability:

The labels that we print for you are available in several formats at

sunsite.unc.edu:/pub/linux/distributions/SLSlabels.

Ordering:

For order information, send mail to DGray@uh.edu. Cost is $6.50 per set and domestic shipping in included. Overseas orders must send payment in US funds and add $1 for shipping. Texas residents add sales tax (6

Entry last modified:

11 Nov 1993.

3.15 Wizvax Communications

Distributor:

Richard Shetron (multics@acm.rpi.edu, multics@wizvax.wizvax.com) Stephanie Gilgut
(stephie@acm.rpi.edu, stephie@wizvax.wizvax.com)

Description:

We sell modem and terminal cables for multi-port serial boards (mainly the Boca Research BB2016, BB1008, BB1004 and the IOAT66). We also sell multi-port serial boards. We can supply other products and services, please ask. We are looking into writting drives for one or more of the intelligent serial boards (probably COMTROL first and perhaps others).

Availability:

     email:  orders@wizvax.wizvax.com  (orders only, please)
             catalog@wizvax.wizvax.com  (automatic response robot)
             multics@acm.rpi.edu or multics@wizvax.wizvax.com

                  stephie@acm.rpi.edu or stephie@wizvax.wizvax.com
                  phone:  (518) 271-6005 (US) 9am-9pm M-sa.
                  Fax:    available soon
                  Snailmail:
                          Wizvax Communications
                          1508 Tibbits Ave.
                          Troy, NY 12180 USA

Ordering:

Send for a catalog (see addresses above). Send orders to address’s listed above. Payment can be personal/company check, US Postal Money Order, Money Order, or Bank check. Orders will not be sent until payment clears. Inquire about PO’s.

Miscellaneous:

wizvax.wizvax.com does a nightly uucp mail exchange. Email may bounce till 05/20/94 as the name servers are being setup.

4 Submissions to this document

Here’s information on how to make submissions to this docuement, and on the policy that we are using for postings to comp.os.linux.announce regarding commercial and mail-order announcements.

4.1 New comp.os.linux.announce policy

One of the purposes of this document is to present short summaries of the many Linux distributions in a compact form for new users to browse over. This document is not merely a concatenation of the various announcements of Linux distributions and the like.

Another purpose of this document is to remove some of the ‘advertisement’ traffic from the newsgroup comp.os.linux.announce. The moderators of c.o.l.a (as well as many of the readers) feel that those people who are re-distributors of Linux distributions which they do not maintain should not be allowed to post periodic advertisements in c.o.l.a.... after all, c.o.l.a is for announcements important to the Linux community; it is not a moderated sales board.

Therefore, if you are distributing Linux software which you do not maintain, the only way that you will be allowed to post information to c.o.l.a is through this HOWTO. Information on submitting entries to this HOWTO is given below. (To submit a posting for c.o.l.a, mail it to linux-announce@tc.cornell.edu. Exceptions are given below.

Here’s an example. If you are, for example, reselling the SLS distribution of Linux on floppies via mail order, the first time that your mail-order service is available, you may post an announcement to c.o.l.a. Thereafter, you must submit an entry in this HOWTO, which will be periodically posted to the newsgroup. Only in the event that you do something dramtically new with your service will you be allowed to post to c.o.l.a about it. (‘Dramatically new’ does not include changing your prices or the versions of the software that you distribute. The moderators have the final say on just what consititutes ‘dramatically new’. :) ) The purpose of this is to reduce the number of periodic postings from people who are using c.o.l.a as an electronic billboard.

However, if you maintain and organize your own distribution of Linux (whether it be available for anonymous FTP, mail order, or both), then you will be allowed to post to c.o.l.a whenever you release a new version. You will not be allowed to make periodic postings about your distribution, in order to ‘advertise’ it every month; only when a new version is released will you be allowed to post to c.o.l.a about it. Of course, you should include information on your distribution in this HOWTO.

I hope that this policy is not too confusing; the one basic rule is that we won’t permit postings to c.o.l.a which are merely advertisements for your distribution or mail order service. The posting must include new information which is of interest to the Linux community at large. Remember that c.o.l.a postings are archived on sunsite.unc.edu; once your posting is out of the news spool it is not gone forever.

4.2 How to submit an entry for this HOWTO

I encourage anyone and everyone who maintains a Linux distribution or mail order service (or consulting service, and the like), to submit information on their service to this HOWTO. It’s easy and fun, and it’s free advertising. This document is posted to many places and is archived (see the next section).

4.2.1 Types of submissions

We are interested in submissions for:

• Complete distributions of Linux software, available either via anonymous FTP, UUCP, or mail order. By ‘complete distribution’ we mean any set of software which can be used to build a complete Linux system from scratch.

• Layered products or individual software packages available only via mail order. If your software package is available via anonymous FTP, chances are people can find it. Software products only available via mail-order include commercial things such as Motif, and any other commercial software ported to Linux.

• Any other LINUX-SPECIFIC goods available via mail order, such as Linux-specific books and documentation, T-shirts, and assorted paraphrenalia.

• LINUX-SPECIFIC services such as telephone or e-mail consulting and technical support. This category also includes miscellany such as development contracts.

If the number of submissions for services and layered products is large, I’ll create a separate HOWTO for these items.

4.2.2 How to submit

To submit an entry to this HOWTO, please send mail to mdw@sunsite.unc.edu with the following information. This ‘format’ is not machine-parsable; any of the fields may be any length that you wish, but I’d like to keep each entry down to, say, 50 lines.

Name:

Name of service or distribution

Distributor:

Name of company, person, etc. who distributes/maintains the service or distribution

Description:

Description of the distribution or service that you provide. If this is a software distribution, please include information such as what software is included, versions, general overview of installation, requirements, and so on.

Availability:

Where your service or distribution is available. This can be an FTP site (including directory pathname, please), a mailing address, phone number, e-mail address, etc.

Ordering:

How to order your distribution or service, if applicable. Include prices, shipping information, methods of payment, etc.

Miscellaneous:

Anything else that you find relevant.

Please keep your entry as short as possible. If you need to include extensive information, please make a reference to where one can FTP or mail to get more information on your distribution; these entries are only meant to be POINTERS to where one can find information on your service or distribution.

If you provide more than one service or distribution, please use SEPARATE ENTRIES for each.

Some things (such as books, t-shirts, etc.) won’t fit exactly into this entry; just be sure to include all relevant information. In other words, this ”entry form” is simply the bare minimum that you must include in your submission to the Distribution-HOWTO; feel free to change, add, or leave out ”fields” as you see fit.

I will more than likely edit your entries to some degree if I find any irrelevant information, or if the entry is overly verbose. Otherwise the content should remain the same.

When making submissions to the Distribution-HOWTO, you grant implicit permission for me to use the entries in other materials, such as books from the LDP, and other online documents. For example, information from the Distribution-HOWTO may be included in a published Linux book. If you do not want me to include your entry in materials other than the Distribution-HOWTO, please say so.

4.3 Acknowledgements

Thanks to Bill Riemers, who produced the first version of the Distribution HOWTO and provided helpful advice. Also, thanks to all of the people who have contributed to this document. I hope that future releases of this HOWTO will be useful to the Linux community; unfortunately, gathering information on the many services out there is a difficult task. Your contributions make this task much easier.